# Searching for periodic signals and quasi-periodic oscillations from an extremely active cycle of FRB 20240114A

### v1

@@ 4.1 Short-Timescale Periodicity in TOAs
A natural expectation from magnetar-based models of repeating FRBs is the presence of underlying periodicities in burst arrival times that trace the rotation of the neutron star. Although a large-scale joint analysis revealed no significant global periodicity, consistent with the non-detection reported by Niu et al. (2022b), we conducted detailed searches within each continuous observation and identified three sets of candidate periodicities, each observed in two separate epochs.
Although we have identified three candidate periodic signals from FRB 20240114A, it is unlikely that all of them correspond to genuine periodicity to the source. Our current selection criteria do not permit us to definitively exclude any of the candidates. While the periods of these candidate signals are on the shorter side of the known magnetar period range, some known magnetars, such as PSR J1846-0258, have even shorter periods of $\sim$0.3266 s (Livingstone et al., 2011).
A striking feature of these candidate signals is the large day-to-day variation in their measured period derivatives ($\dot{P}$), including sign reversals. These variations are difficult to reconcile with the intrinsic spin-down expected from isolated magnetars. Instead, they are more naturally interpreted as apparent period derivatives caused by line-of-sight acceleration in a binary system. If FRB 20240114A resides in such a system, orbital motion could induce variable Doppler shifts, leading to fluctuations in both $P$ and $\dot{P}$ over timescales of days. This scenario is further supported by the fact that the magnitudes of the measured $\dot{P}$ values often exceed those typically associated with magnetar spin-down, sometimes by several orders of magnitude. The system has the potential to yield significant insights into binary dynamics and the role of companion stars in shaping FRB emission.
The absence of substantial periodic detections in our joint and clustered analyses suggests that any underlying periodicity is likely to be transient, intermittent, or strongly modulated, consistent with findings from other FRB studies (Pastor-Marazuela et al., 2023; Du et al., 2025). Du et al. (2025) propose that FRB production may involve multiple independent emission regions in the magnetar magnetosphere, potentially explaining this intermittency. Should these periodicities be confirmed, they would suggest the presence of extremely young, highly magnetized neutron stars as FRB progenitors. If the binary system hypothesis holds, it could also shed light on the role of companion stars in modulating the magnetar’s rotation. Future multi-wavelength observations will be crucial in verifying these periodicities and constraining the system’s properties, including the potential presence of a companion star.
@@ 4.2 Long-Timescale Periodicity in TOAs
In the activity of FRB 20240114A, a candidate periodicity of 69.18 days with a significance of 3.9 $\sigma$ was detected using the epoch folding method, while a similar candidate periodicity of 67.51 days with a significance of 3.5 $\sigma$ was detected using Lomb-Scargle analysis. These results suggest that this repeating FRB may exhibit periodic active behavior around a 68-day period. Concerns about the reliability of this candidate periodicity have been raised due to the uneven distribution of exposure time with respect to phase. Although the Lomb-Scargle periodogram is a useful tool for identifying periodic signals in unevenly sampled data, it is susceptible to observational biases, particularly when the exposure time distribution is uneven. This issue has also been observed in the analysis of other periodic FRBs, such as FRB 20180916B and FRB 20121102A, where the significance of the detected periodicity was influenced by the observation cadence and the exposure time distribution (Chime/Frb Collaboration, 2020; Rajwade et al., 2020).
If the candidate period of approximately 68 days is confirmed, FRB 20240114A would join a growing class of periodic repeating FRBs. The diversity in periods among these sources suggests that multiple possible mechanisms could explain their periodic behavior. For FRB 20180916B, various models have been proposed including neutron star spin periods modulated by disk-wind interactions (Xu et al., 2021), binary orbital modulation with either massive stellar companions or compact objects (Zhang & Gao, 2020; Ioka & Zhang, 2020; Li et al., 2021c), asteroid belt interactions (Dai & Zhong, 2020), and precession mechanisms ranging from free neutron star precession to jet precession in accreting systems (Zanazzi & Lai, 2020; Tong et al., 2020; Chen et al., 2021). Similarly, the $\sim$157-day periodicity of FRB 20121102A has been attributed to either orbital motion or precession effects (Ioka & Zhang, 2020; Levin et al., 2020). The 68-day period of FRB 20240114A, if confirmed, could represent an intermediate case between these established periodic FRBs, potentially favoring binary models with longer orbital periods or different companion types.
The tentative nature of FRB 20240114A’s periodicity highlights the challenges in identifying periodic signals with uneven observational coverage. Future observations should prioritize uniform cadence to minimize sampling biases and multiwavelength follow-up to constrain the local environment. The growing diversity of FRB periods suggests multiple physical mechanisms may be at work, requiring continued observational and theoretical efforts to unravel their origins.
@@ 4.3 Periodicity and QPOs in Burst Time Series
By combining the ACF and the FFT analysis, we reveal temporal features in FRB emissions that are both distinct and complementary. Using the ACF-based method, we identified 4 candidates that exceed the $3\sigma$ threshold. Among them, only one shows four distinct pulses with clearly periodic burst times, indicating a regular temporal spacing. The detected periodicities (ranging from 5 to 62 ms) in the burst clusters exhibit an intriguing behavior: their statistical significance diminishes when the analysis is extended beyond 4 cycles. This suggests a form of transient coherence, possibly originating from unstable oscillation modes in the emission region. These modes may temporarily excite specific periodic components, leading to enhancements in periodicity. The lack of persistent periodic signals over longer timescales is consistent with the stochastic nature of burst arrival times. At the same time, the localized synchronization observed within short windows implies that some degree of phase correlation exists during active emission episodes. Using the FFT analysis, we identified 5 candidates exhibiting statistically significant QPO features. Although the periodogram characteristics could suggest additional candidates, we adopted a conservative selection strategy, retaining only those with unambiguous oscillatory signatures visible in both the time-domain waveforms and their corresponding periodograms.
Figure 7: The result of decomposing the burst time series shown in Figure 5B2 using the EMD. The EMD yields 6 IMFs. Panel a: The original total burst time series. Panel b: The time series obtained by superimposing IMFs 1-3 from the EMD, corresponding to white noise. Panel c: The time series obtained by superimposing IMFs 4-5 from the EMD, corresponding to the QPO time-domain signal. Panel d: The time series obtained from IMF 6 from the EMD, corresponding to the external profile of the burst. Panels e through h: The periodograms corresponding to the time series in panels a through d. The pink curve in panel c represents the best fit of the QPO time-domain profile using model (8).
@@
To further analyze the temporal characteristics of these QPOs, we applied Empirical Mode Decomposition (EMD) to decompose the burst time series into multiple scales. The EMD is an adaptive signal processing technique that decomposes complex nonlinear, non-stationary signals into Intrinsic Mode Functions (IMFs), each representing a specific time-frequency component (Huang et al., 1998; Barbosh et al., 2020). We used a Python implementation of the algorithm to achieve our goal (Laszuk, 2017). We use the QPO candidate shown in Fig. 5 as an example for the analysis. Through the EMD, we obtained 6 IMFs covering different frequency bands. Following the hierarchical transition from high to low frequencies in the IMFs, we combined high-frequency IMFs 1-3 (Component 1), mid-frequency IMFs 4-5 (Component 2), and low-frequency IMF 6 (Component 3), and then calculated periodograms for each component. Figure 7a-d respectively show the original time series and the time series of the three components obtained through the EMD, while Figure 7e-h display their corresponding periodograms. Component 1 primarily consists of high-frequency white noise, and Component 3 represents the envelope of the original burst time series. Component 2 exhibits oscillatory behavior in its time series, with a periodogram consistent with the QPOs in the original periodogram, confirming that Component 2 contains the QPOs extracted via the EMD.
We adopt the method used in Li et al. (2022) to fit the time-domain signals of QPOs, employing the product of a Gaussian function and a sine function to model the QPOs extracted by the EMD. This method is based on the assumption that QPOs are amplitude-modulated signals, and its mathematical expression is
@@
where $A$ represents the maximum amplitude of the signal, $t_{c}$ is the center position of the Gaussian function, $\sigma$ is the width of the Gaussian function, $f$ is the oscillation frequency, and $\phi$ is the initial phase. Using this model to fit the data, we obtain the optimal parameters as $A=1.047$, $t_{c}=0.026$ s, $\sigma=0.003$ s , $f=270.380$ Hz, $\phi=-2.300$ rad. The oscillation frequency $f$ is consistent with the central frequency of the QPOs obtained by directly fitting the periodogram. This further verifies the presence of QPOs in the original burst time series. The corresponding burst time series, QPO time-domain signals extracted via the EMD, the fitting results using equation (8), and their corresponding periodograms are all presented in Appendix C (Figure 9).
Figure 8 displays the distribution characteristics of these QPOs’ central frequencies and full width at half maximum (FWHM). The results show that both the FWHM and the central frequencies of the QPOs exhibit unimodal distributions, with a significant correlation between them (Pearson correlation coefficient $R=0.97$). Linear regression yields a relationship of $y=0.32x+19.99$. This strong correlation suggests that the quality factor ($Q\equiv f_{c}/\mathrm{FWHM}$, where $f_{c}$ is the central frequency of the QPOs) remains relatively stable across different burst time series, indicating that the underlying physical mechanism may be universal.
@@
The microstructure of FRBs provides crucial clues for understanding their radiation mechanisms and the physical properties of their central engines. Among these features, the discovery of QPOs is particularly intriguing. The presence of QPOs may be closely related to various physical processes in compact objects, such as rotation, gravitational lensing, binary star mergers, crustal oscillations in magnetars, or instabilities in the magnetosphere (Piro, 2012; Wang et al., 2016; Wadiasingh & Chirenti, 2020; Pastor-Marazuela et al., 2023). These periodic or quasi-periodic phenomena likely reflect the dynamic behavior of compact objects under extreme magnetic fields or the oscillation modes of magnetospheric plasma. In-depth studies of these QPO features hold the potential to offer new insights into the origin of FRB radiation and their connection to compact objects.
Figure 8: The joint distribution of QPO frequencies and their FWHM in FRB 20240114A, with marginal distributions shown alongside. To avoid harmonic interference, only the lowest frequency QPO from each burst time series is selected as a sample for the plot. The parameters show a strong positive correlation (Pearson $R=0.97$).
@@
A possible physical scenario involves a periodic pulse train superimposed on a broader pulse envelope, which could arise from temporal modulation of the signal strength (e.g., through gravitational lensing effects in pulsars). This mechanism naturally produces a bell-shaped overall profile with short-timescale periodic sub-structures, and may also be considered in the context of FRB emission models. In contrast, the pulse signals we observed do not exhibit strict periodicity, instead showing QPOs. Therefore, the gravitational lensing model is even less effective in explaining the phenomenon we observed. They also explored the possibility of explaining the periodic signal through neutron star mergers. Neutron star mergers are expected to produce QPOs ranging from a few Hz to several kHz, although the merger process is extremely brief. The duration typically depends on the companion star’s mass, ranging from approximately $\sim 10^{2}-10^{4}$ s, or between 0.1 and 1 s (Chime/Frb Collaboration et al., 2022). In contrast, FRB 20240114A is a repeater, which has been erupting for over a year, with no signs suggesting that the bursts will end in the short term. Therefore, the QPOs of FRB 20240114A cannot originate from a neutron star merger.
The microstructure of FRB pulses, especially the appearance of QPOs, may be related to the magnetar’s X-ray burst activity and could be closely associated with the magnetar’s vibrations or intense magnetic field activity (Thompson et al., 2017; Suvorov & Kokkotas, 2019; Li et al., 2021a, 2022).

### v2

@@ 4.1 Short-Timescale Periodicity in TOAs
A natural expectation from magnetar-based models of repeating FRBs is the presence of underlying periodicities in burst arrival times that trace the rotation of the neutron star. Although a large-scale joint analysis revealed no significant global periodicity, consistent with the non-detection reported by the reference [2022RAA....22l4004N], we conducted detailed searches within each continuous observation and identified three sets of candidate periodicities, each observed in two separate epochs.
Although we have identified three candidate periodic signals from FRB 20240114A, it is unlikely that all of them correspond to genuine periodicity to the source. Our current selection criteria do not permit us to definitively exclude any of the candidates. While the periods of these candidate signals are on the shorter side of the known magnetar period range, some known magnetars, such as PSR J1846-0258, have even shorter periods of $\sim$0.3266 s [2011ApJ...730...66L].
The observed magnitude of $\dot{P}$ can be used to estimate possible orbital configurations. Assuming that the entire observed $\dot{f}/f\sim 10^{-7}-10^{-6}\ \mathrm{s^{-1}}$ is due to orbital motion, the corresponding orbital period for a companion mass of $M_{c}=100M_{\odot}$ would be approximately $P_{\rm orb}\sim 0.3-1.9$ days. This range provides a rough constraint on the orbital parameters, indicating that if some of the rapid $\dot{P}$ variations are indeed caused by orbital effects, the orbital period would not be very long.
The absence of substantial periodic detections in our joint and clustered analyses suggests that any underlying periodicity is likely to be transient, intermittent, or strongly modulated, consistent with findings from other FRB studies [2023A&A...678A.149P, 2025arXiv250312013D]. Reference [2025arXiv250312013D] proposes that FRB production may involve multiple independent emission regions in the magnetar magnetosphere, potentially explaining this intermittency. Should these periodicities be confirmed, they would suggest the presence of extremely young, highly magnetized neutron stars as FRB progenitors. If the binary system hypothesis holds, it could also shed light on the role of companion stars in modulating the magnetar’s rotation. Future multi-wavelength observations will be crucial in verifying these periodicities and constraining the system’s properties, including the potential presence of a companion star.
@@ 4.2 Long-Timescale Periodicity in TOAs
Our analysis of FRB 20240114A reveals a significant periodicity at $143.40\pm 7.19$ days with $5.2\sigma$ significance. The coherent phase folding over the entire $\sim$20-month observational baseline demonstrates consistent burst concentration in specific phase regions, regardless of exposure correction. While minor differences exist between exposure-corrected and uncorrected profiles, both show prominent peaks at the same phases, supporting the astrophysical nature of this periodicity. The detection of a harmonically related component at approximately half the fundamental period further supports the astrophysical origin of this signal. The coherence of this periodicity is demonstrated by its consistent alignment with four consecutive active windows. However, the source entered a quiescent state during the final predicted active window. While our observations covered the beginning of this window up to MJD 60949—well within the $\sim$30-day active phase and just 8 days before the predicted peak at MJD 60957—no bursts were detected (see Figure 3A). This demonstrates that while the periodic modulation is a robust and intrinsic property of the source, its expression as detectable bursting is contingent upon the state of a central engine that can transition into quiescence.
With this detection, FRB 20240114A joins the growing class of periodic repeating FRBs, which now exhibits a remarkable diversity in timescales. The population spans from the relatively short 16-day period of FRB 20180916B [2020Natur.582..351C], through the intermediate periods of FRB 20240209A (126 days; [2025ApJ...983L..15P]) and the 143-day period reported here for FRB 20240114A, to the longer $\sim$160-day period of FRB 20121102A [2020ApJ...893L..26I]. This diversity suggests that multiple physical mechanisms may be responsible for the observed periodicity across different sources.
Notably, the 143-day period of FRB 20240114A is remarkably close to the $\sim$160-day period of FRB 20121102A, suggesting these two highly active repeaters may share similar underlying physical mechanisms. Both sources exhibit complex burst morphologies and high activity levels, potentially indicating a common origin for their $\sim$150-day class periodicity. Meanwhile, the much shorter 16-day period of FRB 20180916B likely represents a distinct physical scenario.
For FRB 20240114A’s 143-day period, binary orbital modulation provides a natural explanation. In such models, the period could correspond to orbital motion with either a massive stellar companion or a compact object [2020MNRAS.498L...1Z, 2021ApJ...918L...5L]. The similarity to FRB 20121102A’s period suggests comparable binary separations or system configurations. To assess the consistency of this scenario with other timing features, we compared the expected orbital modulation effects with the properties of candidate short-timescale periodic signals identified in our analysis (Section 4.1). Adopting an extreme companion mass $M_{c}=100M_{\odot}$ for the 143-day orbital period produces a maximum fractional frequency derivative of $\left|\frac{\dot{f}}{f}\right|_{\rm max}\sim 10^{-10}\ \mathrm{s^{-1}}$. This value is several orders of magnitude smaller than the $\dot{f}/f$ variations ($\sim 10^{-7}-10^{-6}\ \mathrm{s^{-1}}$) measured for the candidate short-timescale signals in Table 1. This discrepancy indicates that if the 143-day period represents an orbital modulation, it operates on a distinct timescale from the processes responsible for the observed short-timescale timing variations.
Alternatively, precession mechanisms—including free neutron star precession or jet precession in accreting systems—could produce modulation on such long timescales [2020ApJ...892L..15Z, 2021ApJ...921..147C]. The presence of harmonic structure in FRB 20240114A’s phase profile may provide additional constraints on the system geometry.
Beyond activity periodicity, recent studies have revealed periodic evolution in rotation measures (RM), providing independent evidence for binary system origins. FRB 20201124A shows a 26.24-day RM periodicity [2025arXiv250506006X], while FRB 20220509 exhibits $\sim$200-day RM variations [2025arXiv250510463L]. These RM periodicities strongly support binary system models where the FRB source moves through a magnetized plasma environment modulated by orbital motion.
The detection of this $\sim$143-day period establishes FRB 20240114A as a key member of the periodic FRB population. The existence of multiple period classes ($\sim$16 days, $\sim$126-160 days) suggests that periodicity may be a common feature among repeating FRBs, arising from different physical processes in diverse progenitor systems. The emerging picture from both activity and RM periodicities points toward binary system origins for many repeating FRBs, with orbital characteristics spanning a wide parameter space. Future multiwavelength observations and continued monitoring will be crucial for distinguishing between competing models and understanding the origin of these intriguing periodicities.
@@ 4.3 Periodicity and QPOs in Burst Time Series
By combining the ACF and the FFT analysis, we reveal temporal features in FRB emissions that are both distinct and complementary.
@@
Using the ACF-based method, we identified 4 candidates that exceed the $3\sigma$ threshold. Among them, only one shows four distinct pulses with clearly periodic burst times, indicating a regular temporal spacing. The detected periodicities (ranging from 5 to 62 ms) in the burst clusters exhibit an intriguing behavior: their statistical significance diminishes when the analysis is extended beyond 4 cycles. This suggests a form of transient coherence, possibly originating from unstable oscillation modes in the emission region. These modes may temporarily excite specific periodic components, leading to enhancements in periodicity. The lack of persistent periodic signals over longer timescales is consistent with the stochastic nature of burst arrival times. At the same time, the localized synchronization observed within short windows implies that some degree of phase correlation exists during active emission episodes.
@@
Using the FFT analysis, we identified 2 candidates exhibiting statistically significant QPO features. To further investigate the temporal characteristics of the two detected QPOs, we applied Empirical Mode Decomposition (EMD) to decompose their burst time series into multiple scales. EMD is an adaptive signal processing technique that decomposes complex nonlinear, non-stationary signals into a set of Intrinsic Mode Functions (IMFs), each representing a distinct time-frequency component [1998RSPSA.454..903H, 2020SMaS...29i3001B]. For the No. 1 QPO, the decomposition yielded five IMFs and one residual; for the No. 2 QPO, four IMFs and one residual were obtained. Based on the hierarchical structure of the IMFs from high to low frequency, we grouped IMFs 1–3 (for No. 1) or IMFs 1-2 (for No. 2) as Component 1 (white noise), IMFs 4–5 (for No. 1) or IMF 3-4 (for No. 2) as Component 2 (QPO component), and the residual as Component 3 (representing the low-frequency envelope of the burst). Periodograms were then calculated for each component.
Figures 9a–d in D show the original burst time series and the three components for the No. 1 QPO, while Figures 9e–h display their corresponding periodograms. The same analysis for the No. 2 QPO is shown in Figures 9i–l and 9m–p. Component 1 mainly contains white noise, Component 3 reflects the overall envelope of the burst profile, and Component 2 exhibits clear oscillatory behavior. The periodogram of Component 2 matches the original QPO periodogram, confirming that this component captures the QPO signal extracted through EMD.
To characterize the time-domain behavior of the QPOs, we adopted the fitting method used by the reference [2022ApJ...931...56L], in which the QPO is modeled as the product of a Gaussian envelope and a sinusoidal oscillation. This approach assumes that QPOs are amplitude-modulated signals, described by the following expression:
@@
where $A$ is the peak amplitude, $t_{c}$ is the center of the Gaussian function, $\sigma$ is its width, $f$ is the oscillation frequency, and $\phi$ is the phase. Using this model to fit the data, we obtain the optimal parameters for the No. 1 QPO as $A=1.050$, $t_{c}=0.0260\,\mathrm{s}$, $\sigma=0.0028\,\mathrm{s}$, $f=270.40\,\mathrm{Hz}$, and $\phi=-2.30\,\mathrm{rad}$. For the No. 2 QPO, the fit yields $A=0.534$, $t_{c}=0.0225\,\mathrm{s}$, $\sigma=0.00235\,\mathrm{s}$, $f=324.42\,\mathrm{Hz}$, and $\phi=-3.14\,\mathrm{rad}$. In both cases, the fitted oscillation frequency $f$ is consistent with the peak frequencies obtained from the corresponding periodograms, further confirming the presence of QPOs in the original burst time series.
@@
The microstructure of FRBs provides crucial clues for understanding their radiation mechanisms and the physical properties of their central engines. Among these features, the discovery of QPOs is particularly intriguing. The presence of QPOs may be closely related to various physical processes in compact objects, such as rotation, gravitational lensing, binary star mergers, crustal oscillations in magnetars, or instabilities in the magnetosphere [2012ApJ...755...80P, 2016ApJ...822L...7W, 2020ApJ...903L..38W, 2023A&A...678A.149P]. These periodic or quasi-periodic phenomena likely reflect the dynamic behavior of compact objects under extreme magnetic fields or the oscillation modes of magnetospheric plasma. In-depth studies of these QPO features hold the potential to offer new insights into the origin of FRB radiation and their connection to compact objects.
@@
A possible physical scenario involves a periodic pulse train superimposed on a broader pulse envelope, which could arise from temporal modulation of the signal strength (e.g., through gravitational lensing effects in pulsars). This mechanism naturally produces a bell-shaped overall profile with short-timescale periodic sub-structures, and may also be considered in the context of FRB emission models. In contrast, the pulse signals we observed do not exhibit strict periodicity, instead showing QPOs. Therefore, the gravitational lensing model is even less effective in explaining the phenomenon we observed. They also explored the possibility of explaining the periodic signal through neutron star mergers. Neutron star mergers are expected to produce QPOs ranging from a few Hz to several kHz, although the merger process is extremely brief. The duration typically depends on the companion star’s mass, ranging from approximately $\sim 10^{2}-10^{4}$ s, or between 0.1 and 1 s [2022Natur.607..256C]. In contrast, FRB 20240114A is a repeater, which has been erupting for over a year, with no signs suggesting that the bursts will end in the short term. Therefore, the QPOs of FRB 20240114A cannot originate from a neutron star merger.
The microstructure of FRB pulses, especially the appearance of QPOs, may be related to the magnetar’s X-ray burst activity and could be closely associated with the magnetar’s vibrations or intense magnetic field activity [2017ApJ...841...54T, 2019MNRAS.488.5887S, 2021NatAs...5..378L, 2022ApJ...931...56L]. In this context, recent three-dimensional simulations of magnetar crustal quakes [2025arXiv250812567Q] suggest a plausible origin for the observed QPOs. The $\sim$ kHz seismic oscillations are found to bounce within the crust on millisecond timescales, which may account for the hundreds-of-hertz QPOs detected in our observations and in other FRBs. Furthermore, their simulation demonstrates that such a quake launches a mixture of Alfvén and fast magnetosonic waves into the magnetosphere. The damping of crustal motions on millisecond timescales due to energy loss to the core aligns with the “transient coherence” and short-lived nature of the periodicities we observed. Since the seismic waves are damped before they can spread laterally, the resulting magnetospheric emission remains confined near the quake’s epicenter. This localization may explain why high-significance QPOs appear only intermittently in specific bursts.
The timescale of the microstructure in the radio pulses of pulsar radiation is closely related to the pulsar’s rotation period. Previous studies have shown that for normal, non-recycled pulsars, the quasi-period in their radio pulses is approximately three orders of magnitude smaller than their spin period, i.e., $P_{\mu}\simeq 10^{-3}P$, where $P_{\mu}$ and $P$ are the quasi-period and the spin period of the pulsar’s radio pulses, respectively. Since we detected some QPOs in FRB 20240114A, if it indeed originates from a magnetar, we can attempt to estimate the magnetar’s period based on the known relationship between the quasi-period of substructures and the spin period. The relationship is given by $P_{\mu}=(0.94\pm 0.04)\times P^{0.97\pm 0.05}\,\text{ms}$, which has been confirmed in both normal pulsars and magnetars and is applicable over about six orders of magnitude [2024NatAs...8..230K]. Following the method in the reference [2024NatAs...8..230K], we computed the geometric mean of all detected quasi-periods and estimated the magnetar’s spin period to be in the range of approximately 3 s to 76 s.
The detection of QPOs in only a handful of bursts among tens of thousands cannot be easily explained by sample size or signal-to-noise considerations alone. This sporadic appearance aligns with previous observations that QPOs are not always present, even under similar conditions. For instance, the magnetar PSR J1622–4950 displayed clear QPO-like structures in 2017 observations, yet they were much weaker or absent in data from early and late 2018 [2024NatAs...8..230K], reinforcing the notion that their occurrence is inherently variable rather than purely observational in origin.
5 Summary
This study presents a multi-timescale analysis of (quasi-)periodic signals in the repeating fast radio burst FRB 20240114A. The key findings are:
  1. 1.